OpenStudy (anonymous):

#please help me find the two positive numbers that satisfy the given requirements. 1. the sum is S and the product is a maximum

OpenStudy (phi):

is this a calculus problem?

OpenStudy (dape):

If their sum, S, is even, choose the two numbers to be S/2, since a square maximizes the area (the product of the two numbers) with a given perimeter. If their sum is odd, then it can be written S=2n+1=n+(n+1), so let one number be n and the other n+1, since that's the closest to a square.

OpenStudy (anonymous):

yes

OpenStudy (phi):

let the 2 numbers be x and y x+y= S y= S-x find the max of x*y = max of x(S-x) You have a function f(x)= x(S-x) to find its max, take the derivative, set = 0, and solve for x

OpenStudy (dape):

Oh, I misread, thought it was only for integers. Just let the smaller number be a and the bigger b, then you can write b as b=S-a. And use that to minimize the product P=a*b=a*(S-a).

OpenStudy (phi):

**to find its max to show you have a max (rather than a min),take the 2nd derivative and show it is negative.

OpenStudy (anonymous):

how can i show it is negative? can you give the 1st and 2nd derivative?

OpenStudy (phi):

\[ f(x)=x(S-x) = Sx-x^2\] where S is constant. Can you find the derivative?

OpenStudy (anonymous):

S-2x

OpenStudy (anonymous):

?

OpenStudy (phi):

yes. and the 2nd derivative is -2, which means you will be finding a maximum when you solve S-2x=0

OpenStudy (anonymous):

what is the maximum?

OpenStudy (phi):

you solve for x in S-2x=0 you found the derivative of f(x). this is the slope. when the slope is 0, you are at a max